typical heights and wieghts for adult men are in the table below
height (in inches) 70 72 74 weight (in pounds) 170 180 190

explain why these data are linear?

Answer by checkley77(12569)

HERE WE NEED TO FIND THE SLOPES BETWEEN ALL POINTS.
(180-170)/(72-70)=10/2=5
(190-170)/(74-70)=20/4=5
(190-180)/(74-72)=10/2=5
BECAUSE THE SLOPE IS 5 FOR ALL LINES THEN THE 3 POINTS ARE ON THE SAME LINE THUS THESE 3 POINTS FORM A LINEAR LINE.